If Steven can mix 20 drinks in 5 minutes, Sue can mix 20 drinks in 10 minutes, and Jack can mix 20 drinks in 15 minutes, how much time will it take all 3 of them working together to mix the 20 drinks?

Answer by jorel1380(3719)

20/5+20/10+20/15=20/x
120x+60x+40x=600
220x=600
x=600/220=30/11 mins..